OpenStudy (anonymous):

lim [abs(x-7)] / [x-7] x -> 7 (-)

OpenStudy (anonymous):

how do you deal with the (-)? i think that means from the left? right? how do you find the limit form only one direction?

OpenStudy (anonymous):

and how does that compare to: lim [abs(x-7)] / [x-7] x -> 7 where you find the limit form both directions?

OpenStudy (anonymous):

(-) means take its limit from the left

OpenStudy (anonymous):

and this means that you should multiply the absolute value by -1

OpenStudy (anonymous):

7-x/x-7=-1

OpenStudy (anonymous):

so, is the limit from the left -1?

OpenStudy (anonymous):

but the limit from both sides is positive 1? right?

OpenStudy (anonymous):

yes, it would be +1 from the right which means limit does not exist at all, for limit from the left is not equal to the limit from the right

OpenStudy (anonymous):

'cause you mulitply it by 1?

OpenStudy (anonymous):

yes

OpenStudy (anonymous):

aha!

OpenStudy (anonymous):

in fact the point is not to multiply it by 1 but to understand the absoulte value will be the same if the limit is taken from the right side

OpenStudy (anonymous):

I have one more question...

OpenStudy (anonymous):

it's about the intermediate value theorem

OpenStudy (anonymous):

f(x) = x^3*cos(x) - x*sin(x) + 1 in the interval [-2,2] they want me to prove that there are at least two zeros within the interval... what does that mean? like, where they cross the axis? do you plug in zeros for the x? how do you find more than one zero? there's only one solution if you plu in the zero?

OpenStudy (anonymous):

or, should i post this as a question?

OpenStudy (anonymous):

please let me check for a second

OpenStudy (dumbcow):

plug in the x_values -pi/2, 0, pi/2 f(-pi/2) = 0 - (-pi/2)(-1) +1 = -.57 f(0) = 0 -0+1 = 1 f(pi/2) = 0 -(pi/2)(1) +1 = -.57 f(x) goes from - to + to -, since it is a continuous function it proves f(x) = 0 at least twice in that interval